OpenStudy (anonymous):

Find f. f ''(x) = 8 − 18x, f(0) = 7, f(2) = 11

OpenStudy (perl):

you can integrate that twice

OpenStudy (anonymous):

f''(x) = 4x^2 − 3x^3 +C

OpenStudy (perl):

you should get two constants

OpenStudy (anonymous):

like this? f''(x) = 4x^2 − 3x^3 +Dx + C

OpenStudy (perl):

$$\Large { f ' ' (x) = 8 - 18x, \\ \\ \therefore \\ f (x) = 4x^2 - 3x^3+ C_1x + C_2\\ f(0) = 4(0)^2 - 3(0)^3+ C_1(0) + C_2 = 7 \\ \\ \therefore \\ C_2 = 7 } $$